A sample of n=5 scores has a mean x-bar=8.If a new score X=2 is added to the sample.What value would be obtained for the new sample?
1
Expert's answer
2012-01-06T07:27:02-0500
We know that (a1+a2+a3+a4+a5)/5 = 8. So, (a1+a2+a3+a4+a5+2)/6 = (a1+a2+a3+a4+a5)/6 + 2/6 = ((a1+a2+a3+a4+a5)/5)*5/6 + 2/6 = 8*5/6 + 1/3 = 7.